# Safe Flow Matching: Robot Motion Planning with Control Barrier Functions

### v1

@@ 4.1 Safe Flow Matching under Unified Constraints
s.t. $\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ $\forall i\in\{0,\cdots,H\}.$ (9b)
Given that $\|\bm{u}\|^{2}=\sum_{i=0}^{H}\|\bm{u}^{i}\|^{2}$, the optimization problem can be decoupled for each $\bm{u}^{i}$ as
$\bm{u}_{t}^{i}=\operatorname*{arg\,min}_{\bm{u}^{i}\in\mathbb{R}^{d_{s}}}\|\bm{u}^{i}\|^{2}$ (10a)
s.t. $\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ (10b)
which can be consider as a quadratic programming (QP) problem due to the quadratic objective function and linear constraint w.r.t $\bm{u}^{i}$. To alleviate the derivation of the solution of this QP problem, we first introduce two auxiliary variables as follows
$\bm{b}_{t}^{i}=\left(\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\right)^{\top},$ (11)
$a_{t}^{i}={\bm{b}_{t}^{i}}^{\top}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t}).$ (12)
By constructing a valid FMBF to ensure $a_{t}^{i}\geq 0$ when $\bm{b}_{t}^{i}=\bm{0}_{d_{s}\times 1}$, then the solution for such a QP problem always exists. Specifically, the explicit solution is written as
$\bm{u}_{t}^{i}=\begin{cases}\bm{0}_{d_{s}\times 1},&\text{if~{}}a_{t}^{i}\geq 0\\ -{\bm{b}_{t}^{i}a_{t}^{i}}/{\|{\bm{b}_{t}^{i}}\|^{2}},&\text{otherwise}\end{cases}$ (13)
for all $i=0,\cdots,H$ and $t\in[0,1)$. Note that the function $\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))$ is not defined at $t=1$. Therefore, the determination method for $\bm{u}_{t}$ with $t\in[0,1)$ is not implementable. Considering $\mathcal{T}_{1}$ plays the most important role as output, For $t=1$ with $\mathcal{T}_{1}=\bm{\psi}_{1}(\mathcal{T}_{0})$, a safety filter is directly designed for $\mathcal{T}_{1}$ as
[…]
which directly ensure the trajectory safety for $\mathcal{T}_{1}$.
@@ 4.2 Safe Flow Matching under Composite Constraints
Defining a safe set using a single FMBF may be inadequate for capturing intricate or multiple safety constraints. In this subsection, we introduce a composite safe defined by $N$ differentiable continuous functions $h_{j}(\cdot):\mathbb{S}\to\mathbb{R}$ with $j=1,\cdots,N$. Then a composite safe set $\mathbb{C}_{s}$ is defined as
$\mathbb{C}_{s}=\bigcap_{j=1}^{N}\mathbb{C}_{s,j},$ $\mathbb{C}_{s,j}=\{\bm{s}\in\mathbb{S}:h_{j}(\bm{s})\geq 0\},$ $\forall j=1,\cdots,N$ (15)
We extend the definition of trajectory safety, safeFM, and FMBF from the unified case to this composite framework, as detailed below.
@@
The function $\bm{h}(\cdot)$ is called a composite flow matching barrier function (CFMBF), if it satisfies the following conditions:
@@
    hold for all $j=1,\cdots,N$, where
    $\varphi_{j}(t,h_{j}(\bm{s}))=\begin{cases}\varphi_{0,j},&\text{if~{}}h_{j}(\bm{s})\geq 0\\ \varphi_{1,j}(t),&\text{otherwise}\end{cases}$ (17)
    with $\varphi_{0,j}\in\mathbb{R}_{+}$ and a blow-up function $\varphi_{1,j}(\cdot):[0,1)\to\mathbb{R}_{+}$ satisfying $\lim_{t\to 1^{-}}\varphi_{1,j}(t)=+\infty$.
@@
Definition 7 (Trajectory Safety with Composite FMBF).
A trajectory $\mathcal{T}$ of length $H$ is safe if satisfies the condition in Definition 3 w.r.t $h_{j}(\cdot)$ for all $i=1,\cdots,N$.
Definition 8 (Safe Flow Matching with Composite FMBF).
A flow matching process is called safe w.r.t $\bm{h}(\cdot)$, if the generated trajectory $\mathcal{T}_{1}$ at $t=1$ is safe w.r.t $\bm{h}(\cdot)$ defined by Definition 7 for any initial sample $\mathcal{T}_{0}\sim p_{0}$.
Building on the safety guarantees established for a unified safe set, we demonstrate that a composite flow matching barrier function ensures the safety of both the trajectory and the flow matching process in the composite case, as formalized below.
@@
Same as the seriaro in the unified contracts, we incorporate a minimal regularization term $\bm{u}_{t}$ into the flow matching process while ensuring safety under composite constraints. For SafeFM under Composite Constraints, the QP problem is formulated to determine $\bm{u}_{t}$ for any $t\in[0,1)$ as
@@
Then the problem can be decoupled into independent subproblems for each $\bm{u}_{t}^{i}$, such that
$\bm{u}_{t}^{i}=\operatorname*{arg\,min}_{\bm{u}^{i}\in\mathbb{R}^{d_{s}}}\|\bm{u}^{i}\|^{2}$ (19)
s.t. $\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ $\forall j\in\{1,\cdots,N\}$
However, with multiple linear constraints, the feasibility of this decoupled QP is not guaranteed. To ensure feasibility, we introduce relaxation terms $\delta_{j}^{i}\in\mathbb{R}_{0,+}$ and reformulate decoupled QP problem as
$\bm{u}_{t}^{i}=\operatorname*{arg\,min}_{\bm{u}^{i}\in\mathbb{R}^{d_{s}},\delta_{j}^{i}\in\mathbb{R}_{0,+}}\|\bm{u}^{i}\|^{2}+\sum_{j=1}^{N}{\delta_{j}^{i}}^{2}$ (20a)
s.t. $\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\mathrm{d}h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\mathrm{d}\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})+\delta_{j}^{i}\geq 0,$ (20b)
$\forall j\in\{1,\cdots,N\}.$
Remark 1.
It is straightforward to verify that choosing sufficiently large values for $\delta_{j}^{i}$ always ensures the feasibility of the QP problem. This effectively turns the problem into a feasibility problem. For instance, one feasible solution is $\bm{u}_{i}=\bm{0}_{d_{s}\times 1}$ and $\delta_{j}^{i}=\max\{0,-\frac{\partial h_{j}(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})-\varphi_{j}(t,h_{j}(\bm{E}_{i}\mathcal{T}_{t}))h_{j}(\bm{E}_{i}\mathcal{T}_{t})\}$ for all $j=1,\cdots,N$.
Remark 2.
Due to such relaxation for feasibility, the positivity for $h_{j}(\bm{E}_{i}\mathcal{T}_{1^{-}})$ may not be satisfied. In such cases, the importance of condition 2 in Definition 6 is shown, where the safe trajectory $\mathcal{T}_{1}$ is obtained by solving
[…]
If $\mathbb{C}_{s}$ is non-empty, then (21) is feasible. The non-empty $\mathbb{C}_{s}$ indicates the existence of $\bm{s}_{safe}\in\mathbb{C}_{s}\subseteq\mathbb{S}$. Then, the trajectory $\mathcal{T}_{safe}=[\bm{s}_{safe}^{\top},\cdots,\bm{s}_{safe}]^{\top}$ is safe as in Definition 8, which indicates the existence of at least one solution, i.e., $\mathcal{T}_{1}=\mathcal{T}_{safe}$, for (21).
@@ 5 Experimental Results
5.1 Planar Robot Navigation
We demonstrate the effectiveness of SafeFM in a planar robot navigation scenario. We generate category-specific trajectories within a two-dimensional environment, verifying the quality and safety of generated trajectories.
Task Definition.
This experiment evaluates the performance of our SafeFM model for planar robot navigation. Specifically, the model generates two-dimensional trajectories by numerically integrating a learned deterministic vector field from random initial points sampled from a simple prior. Trajectories are categorized into two distinct types determined, determined by navigation directions from the lower-left to upper-right and from the upper-left to lower-right, respectively. The objective is twofold:
  • Verify whether the generated trajectories start and terminate within predefined regions.
  • Assess the effectiveness of safeFM in regulating trajectories to avoid obstacles.
Data and Model.
The FM model was trained on a synthetic dataset comprising two classes of planar trajectories, each with $10,000$ samples. The FM model was trained on a synthetic dataset consisting of two classes of planar trajectories, each with 10,000 samples. Every trajectory is represented as a 100-point Bézier curve, with its start and end points sampled uniformly from circles of radius 0.2. For class 0, the start and end circles are centered at (-1,-1) and (1,1), respectively; for class 1, they are centered at (-1,1) and (1,-1). During training, the model learned to map random noise to smooth, goal-directed trajectories following specific directional distributions. At inference, the model parameters remain fixed, and trajectories are generated by numerically solving an ODE initialized from random points. The CBF module subsequently corrects trajectories to enforce safety constraints.
Evaluation Metrics.
The following metrics are used to comprehensively evaluate trajectory quality and safety:
  • Success Rate: Proportion of generated trajectories with start and end points within predefined goal regions.
  • Safe Trajectory Ratio: Proportion of trajectory waypoints satisfying safety constraints, providing further insight into the effectiveness of the CBF projection.
  • Inference Efficiency: Average computational time required to generate trajectories, indicating the model’s suitability for real-time applications.
Results and Analysis.
To illustrate the effectiveness of SafeFM in generating smooth, safe, and goal-directed trajectories for planar robot navigation, we qualitatively analyze the trajectory generation process based on benchmark experiments. We conducted extensive evaluations using $1,000$ trajectory samples per class to assess both the accuracy and safety of the generated trajectories. Starting from random initial noise patterns, SafeFM model consistently and smoothly converges to trajectories that accurately reflect the learned class-specific characteristics. These generated trajectories clearly capture the directional patterns and spatial distributions inherent in the original training dataset, demonstrating the model’s ability to learn and generalize critical trajectory features for each navigation scenario.
To further confirm endpoint accuracy, Figure 1 visualizes representative trajectories from each class, along with predefined circular regions (radius = 0.2) marking the desired start and end zones. All displayed trajectories consistently initiate and terminate within these designated goal regions, demonstrating that the FM model reliably adheres to endpoint constraints.
Figure 1: Representative trajectories with goal regions (radius = 0.2). Trajectories consistently start and end within the designated zones.
Figure 2 highlights the critical role of the CBF in trajectory safety correction.
Figure 2: Trajectory safety improvement via CBF projection. Obstacle region (radius = 0.25) is centered at the origin. Trajectories after projection avoid unsafe regions.
As shown in Table 1, the baseline FM model achieves reasonably high accuracy at both start and end positions, with endpoint success rates exceeding 86% for both classes. However, FM model generated trajectories also exhibit a notable percentage of obstacle violations 12.40% for Class 0 and 12.73% for Class 1 indicating safety concerns in cluttered environments. The SafeFM model eliminates all obstacle violations entirely across both classes. In addition, SafeFM improves both start and endpoint accuracies to 100.00%, demonstrating its capability to simultaneously enforce trajectory safety and enhance positional precision. The detailed results are summarized in Table 1.
Table 1: Trajectory Accuracy and Safety Evaluation
Method Class Start Accuracy (%) End Accuracy (%) Obstacle Violation (%)
FM 0 88.60 86.40 12.40
1 86.80 90.10 12.73
SafeFM 0 100.00 100.00 0.00
1 100.00 100.00 0.00
Overall, these quantitative results clearly demonstrate the effectiveness of SafeFM in generating accurate and collision-free trajectories, highlighting its potential for real-world deployment in planar robotic navigation scenarios.

### v2

@@ 4.1 Safe Flow Matching under Unified Constraints
s.t. $\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{E}_{i}\bm{v}_{t}^{\theta}(\bm{\tau})+\frac{\partial h(\bm{E}_{i}\mathcal{T}_{t})}{\partial\bm{E}_{i}\mathcal{T}_{t}}\bm{u}^{i}+\varphi(t,h(\bm{E}_{i}\mathcal{T}_{t}))h(\bm{E}_{i}\mathcal{T}_{t})\geq 0,$ $\forall i\in\{0,\cdots,H\},$ (8b)
which can be considered as a quadratic programming (QP) problem due to the quadratic objective function and linear constraint w.r.t $\bm{u}_{t}$. The explicit closed-form expression of $\bm{u}_{t}$ refers to Appendix B.
@@ 4.2 Safe Flow Matching under Composite Constraints
For a complex safe set, defining a safe set using a single FMBF may be inadequate for capturing intricate or multiple safety constraints. In this subsection, we introduce a composite safe set defined by $N$ differentiable continuous functions $h_{j}(\cdot):\mathbb{S}\to\mathbb{R}$ with $j=1,\cdots,N$ and $N\in\mathbb{N}_{>0}$. Then a composite safe set $\mathbb{C}_{s}$ is defined as
$\mathbb{C}_{s}=\bigcap_{j=1}^{N}\mathbb{C}_{s,j},$ $\mathbb{C}_{s,j}=\{\bm{s}\in\mathbb{S}:h_{j}(\bm{s})\geq 0\},$ $\forall j=1,\cdots,N.$ (9)
We extend the definition of trajectory safety, safe flow matching, and FMBF from the unified case to this composite framework, as detailed below.
@@
The function $\bm{h}(\cdot)=[h_{j}(\cdot)]_{j=1,\cdots,N}$ is called a composite flow matching barrier function (CFMBF) if it satisfies the following conditions:
@@
    holds for all $j=1,\cdots,N$, where $\varphi_{j}(t,h_{j}(\bm{s}))=\varphi_{0,j}$ if $h_{j}(\bm{s})\geq 0$ with $\varphi_{0,j}\in\mathbb{R}_{>0}$, and $\varphi_{j}(t,h_{j}(\bm{s}))=\varphi_{1,j}(t)$, otherwise. The blow-up function $\varphi_{1,j}(\cdot):[0,1)\to\mathbb{R}_{>0}$ monotonically increases satisfying $\lim_{t\to 1^{-}}\varphi_{1,j}(t)=+\infty$ and $\int_{0}^{1^{-}}\varphi_{1,j}(s)\mathrm{d}s\!=\!\infty$.
@@
Definition 7 (Trajectory Safety with CFMBF).
A trajectory $\mathcal{T}$ of length $H+1$ is safe if satisfies the condition in Definition 3 w.r.t $h_{j}(\cdot)$ for all $j=1,\cdots,N$.
Definition 8 (Safe Flow Matching with CFMBF).
A flow matching process is called safe w.r.t. composite $\bm{h}(\cdot)$, if the generated trajectory $\mathcal{T}_{1}$ at $t=1$ is safe w.r.t. composite $\bm{h}(\cdot)$ defined by Definition 7 for any initial sample $\mathcal{T}_{0}\sim p_{0}$.
Building on the safety guarantees established for a unified safe set, we demonstrate that a CFMBF ensures the safety of both the trajectory and the FM process in the composite case, as formalized below.
@@
Same as the scenario in the unified contracts, we incorporate a minimal regularization term $\bm{u}_{t}$ into the flow matching process while ensuring safety under composite constraints. For SafeFlow under composite constraints, the QP problem is formulated to determine $\bm{u}_{t}$ for any $t\in[0,1)$ as
@@ 5 Experimental Results
We conduct simulation experiments to answer the following questions:
  • Q1:
    Does SafeFlow generate obstacle-avoiding trajectories with provable safety guarantees?
  • Q2:
    What impact do FMBFs have on the diffusion and flow-based model in the inference process?
  • Q3:
    How does SafeFlow perform in comparison to SOTA methods that incorporate safety constraints?
In particular, we compare against several SOTA methods, including all variants of SafeDiffuser [22]: Robust-Safe Diffuser (RoSD), Relaxed-Safe Diffuser (ReSD), and Time-Varying-Safe Diffuser (TVSD), as well as CoB and CoBL from [23], which combine control barrier functions with or without Lyapunov functions within the diffusion framework. Besides the conventional guidance-based approaches, such as Classifier Guidance (CG) [29] and Truncate (TRUNC) [30], we implement our proposed FMBFs on top of the diffusion model, named SafeDiffus, to ensure safety guarantees.
5.1 Robot Navigation
In this subsection, we consider a simple task that the robot can navigate in planar directions from the lower-left to upper-right as illustrated in Fig. 2. This experiment evaluates the performance of SafeFlow and SafeDiffus and other methods for planar robot navigation with six metrics [^2] Additional details and results are provided in Appendix C..
  • The safety rate (Safety) indicates the proportion of generated trajectories that avoid collisions with obstacles.
  • The Kullback–Leibler (KL) divergence of goal state distributions (KL-G) is utilized to quantify the discrepancy between the generated data distribution and the true distribution observed in the dataset, which indicates the reachability of the generated final states.
  • Moreover, we use curvature-based smoothness (CS) to quantify the smoothness of discrete trajectories, which captures the degree of directional change between consecutive path segments, and acceleration-based smoothness (AS) to assess the change in velocity.
  • Inference time (Time) measured the duration required to generate a single trajectory.
To illustrate the effectiveness of SafeFlow in generating safe, smooth, and goal-directed trajectories for planar robot navigation, we qualitatively analyze the trajectory generation process based on benchmark experiments with $1000$ trials in Table 1. Notably, SafeFlow achieves a perfect safety rate of 100%, significantly outperforming all baselines with safety guarantees, which provides both experimental and theoretical support for Q1. It is noted that the safety claimed in [23, 22] is only probabilistic due to the stochastic property of the diffusion model and the utilization of conventional CBF, which induces low safety in our scenarios. To answer Q2, we further analyze the differences between the vanilla FM and SafeFlow, as well as the vanilla Diffuser and SafeDiffus. SafeFlow and SafeDiffus incur approximately 60% higher inference time due to the safety regularization. Nevertheless, they outperform other baselines that incorporate safety strategies. The increases in KL-G and CS are expected, as the models need to make directional adjustments to avoid obstacles. Notably, the lower CS values of FM and the vanilla Diffuser reflect overly simplistic, straight-line trajectories that compromise safety. Interestingly, in the AS metric, our SafeDiffus outperforms the vanilla Diffuser, while FM and SafeFlow and FM yield the lowest AS scores. The answer to Q3 is illustrated in Fig. 2, which visually compares the trajectory generation performance across different methods. It is evident that SafeFlow effectively avoids obstacles while closely adhering to the trajectory distribution observed in the training dataset. SafeDiffus also achieves zero collisions; however, some outliers remain, where a few generated trajectories exhibit higher sensitivity. In contrast, other approaches either collide with obstacles or generate trajectories that deviate significantly from the training distribution. This highlights the ability of the proposed framework to outperform SOTA methods, offering both training-free deployment and test-time safety guarantees.
Figure 2: Comparison of generated trajectories from 15 samples across different methods.
Table 1: Safe planning for robot navigation comparisons with benchmarks.
Method Safety $\uparrow$ KL-G $\downarrow$ CS $\downarrow$ AS $\downarrow$ Time (s) $\downarrow$
CG [29] 10.39% 0.0252 0.2492 0.0152 0.96
TRUNC [30] 11.24% 0.0254 0.4668 0.6131 0.93
RoSD [22] 5.42% 0.0303 0.4747 0.0837 29.84
TVSD [22] 15.51% 0.0241 0.3283 0.0311 27.62
ReSD [22] 9.38% 0.0257 0.0629 0.1068 32.54
CoB [23] 9.59% 0.0248 0.5548 0.6561 1.29
CoBL [23] 2.40% $\mathrm{>10^{16}}$ 0.5102 0.7294 4.32
Diffuser [2] 10.58% 0.0224 0.0522 0.2858 0.31
SafeDiffus (Ours) 100.00% 0.0233 0.0985 0.0589 0.54
FM [7] 10.96% 0.0265 0.0040 0.0001 0.26
SafeFlow (Ours) 100.00% 0.0291 0.1150 0.0062 0.38
5.2 Robot Manipulation
To demonstrate the effectiveness of the proposed SafeFlow and SafeDiffus methods, we evaluate the same metrics used in Section 5.1 for a robotic manipulation task using Franka Research 3as shown in Fig. 3(a). Specifically, the objective is to move the robot arm from a lower to an upper position without collision with obstacles, which are unseen in the training data sets. The experiment is repeated $1000$ times for each method, and the performance of each method is evaluated based on the metrics mentioned in Section 5.1 for the motion trajectory of the end-effector, which is summarized in Table 2. It is obvious to see that the proposed SafeFlow demonstrates clear superiority in KL-G, CS, and inference time (except vanilla diffusion and flow models), especially in the safety aspect with 100% safety guarantees, providing a strong affirmative to Q1. It is noted that some baseline methods (Diffuser, FM, and ReSD) have low CS or AS due to the generation of overly simplified trajectories, lacking the effective response to obstacles and inducing insecurity. Therefore, SafeFlow demonstrates better performance compared to the SOTA methods [22, 23, 29, 30] that integrate safe constraints, which answers Q3. A similar phenomenon observed in Section 5.1 is that there is a slight degradation in performance compared to vanilla approaches (Diffsuer and FM) is expected, as the FMBFs influence the generation process during inference to enforce safety.
(a) Franka Research 3
(b) CoB
(c) TVSD
(d) SafeFlow (Ours)
Figure 3: End-effector trajectories for selected methods. (a) Tested Franka Research 3. (b-c) Demonstration of end-effector trajectories from experiments, where (b,c) are chosen with the highest safety rate in [22, 23].
Table 2: Safe planning for robot manipulation comparisons with benchmarks.
Method Safety $\uparrow$ KL-G $\downarrow$ CS $\downarrow$ AS $\downarrow$ Time $\downarrow$
CG [29] 27.63% 0.3221 0.6897 0.0223 20.02
TRUNC [30] 33.31% 0.1199 0.6734 0.0135 19.80
RoSD [22] 68.22% 0.2246 0.6168 0.0227 40.16
TVSD [22] 69.31% 0.1219 0.6855 0.0253 123.47
ReSD [22] 39.28% 0.1436 0.6729 0.0129 66.46
CoB [23] 34.51% 0.4257 0.6528 0.0345 20.16
CoBL [23] 23.21% 0.5668 0.6330 0.0497 20.29
Diffuser [2] 10.03% 0.0378 0.6207 0.0126 16.61
SafeDiffus (Ours) 100.00% 0.0711 0.6542 0.0408 41.65
FM [7] 13.98% 0.0227 0.1006 0.0196 5.13
SafeFlow (Ours) 100.00% 0.0612 0.1734 0.0251 14.76